OpenStudy (anonymous):

Raman's present age is three times his daughter's and nine-thirteenth of his mother's present age. The sum of the present ages of all three of them is 125 years. What is the difference between the present ages of Raman's daughter and Raman's mother?

OpenStudy (anonymous):

Kamya purchased an item for 746,000 and sold it at a loss of 12 per cent. With that amount she purchased another item and sold it at a gain of 12 per cent. What was her overall gain/loss?

OpenStudy (mertsj):

r = raman d=daughter m=mother r = 3d, so d = 1/3r r=9/13m so m = 13/9r r+1/3r+13/9r=125 Multiply by 9 and solve